Week of 160418

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:
  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Luca (SCOD+Storage), Edoardo (Network), Andrea M. (FTS), Maria D. (GGUS), Miguel (Compute)
  • remote: Dario (ATLAS), Stefano (CMS), Marcelo (LHCb), Michael (BNL), Antonio (CNAF), Rolf (IN2P3), Andrew (Nikhef), Dimitri (KIT), Dmytro (NDGF), Kyle (OSG), John (RAL), Di Qing (TRIUMF), Pepe (PIC - not able to connect)

Experiments round table:

  • ATLAS reports (raw view) -
    • Activities:
      • MC15c (digi+reco reconstruction) ongoing
      • Heavy Ion reprocessing almost completed. Memory consumption reduced through less monitoring histos and a few other fixes.
      • Online->EOS->Castor throughput test ran OK.
      • All sites associated to WORLD cloud last Friday. Now Tier-3s and non-MCP Tier-2s used again for production (and full).
      • On-going discussion on queue attributes needed for job matching (especially memory).
    • Problems:
      • Discussion ongoing on working directory definition for ACT+ARC-CE jobs running on TW sites.
      • Changes to FTS job submission last Thursday produced an increase of jobs queued in FTS but not polled for status. To be discussed by the Rucio and FTS people.

  • CMS reports (raw view) -
    • Spring16DR80 production ongoing. All resources used. Multicore deployment at T2 progressing.
    • No problem to report

  • ALICE -
    • Apologies: operations experts are at the ALICE T1-T2 workshop in Bergen.
    • VOMS-Admin not working this morning: GGUS:120890
      • it turned out the client side networks were filtering non-standard ports like 8443,
        i.e. nothing wrong on the server side
    • intermittent expired certificate errors for GGUS web pages: GGUS:120891

  • LHCb reports (raw view) -
    • Activity
      • Staging activities on all T1 sites for 2012 data stripping
    • Site Issues
      • T0: Problem uploading to EOS Disk , fixed on 12th (GGUS:120799)
      • T1:
        • SARA FTS problems when staging in data (GGUS:120808), fixed
        • IN2P3 gridftp problem to access files (GGUS:120835), fixed
        • RAL network problem is fixed
        • RAL disk server is still being investigated (some files inaccessible)
        • GRIDKA running few jobs is fixed (increased priority for the jobs) (GGUS:120874), fixed
        • GRIDKA: GGUS was unstable this morning, back already

Sites / Services round table:

  • ASGC:
  • BNL: NTR
  • CNAF: LHCb disk is full. We should solve in a week, after adding around 500TB to the storage. Currently we are a bit under the pledge with this addition we should provide it.
  • FNAL:
  • GridPP:
  • IN2P3: NTR
  • JINR:
  • KISTI: NTR (Sang is attending the ALICE Tier workshop at Bergen)
  • KIT:
    • Downtime next week (Wed and Thu) for the tape infrastructure affecting all experiments
    • Ticket to exclude service for availability count GGUS:120591
  • NDGF:
    • Frequent issues in connecting with VOMS
    • Problem with the CERN CA generated by the OCSP,the firewall port on the machine seems closed, currently the check has been disabled to avoid further issues
Luca suggested to open GGUS ticket to keep track of these issues, in particular the VOMS related one. Andrea M. said that a similar issue was already reported in GGUS:120784. Maria D. and Andrea M. will report about this in the next WLCG Operation Coordination meeting.

  • NL-T1: NTR
  • NRC-KI:
  • OSG: investigation ongoing for an issue in the bridge between GGUS and the OSG ticketing system
  • PIC: NTR
  • RAL: There were network problems at RAL over the weekend, from midnight on Friday until 14:30 BST on Saturday. A core network switch had an issue and it caused repeated interruptions to service. The networks team had to attend site to fix it.
  • TRIUMF:

  • CERN Network:
    • CERN-BNL: transparent maintenance on the main link on Wed morning, traffic will be redirected on the second link
    • CERN-IN2P3: second 10Gb link in production, the traffic is now loadbalanced on two links
  • CERN computing services:
    • Lost connectivity to Wigner today (18th April) around 11am, LSF and Condor affected OTG:0029980 OTG:0029984
    • On 14th of March, SAM production instances were migrated to the new hypervisors, which had an impact on the performance of the VMs they are running on and caused the instability observed and reported by the sites and experiments. The issue has been investigated and a regular disk arrays check has been identified as the main source of the high IO wait observed during the weekends, it was disabled on Wed last week (13th of April). A general re-computation was done for March and will be done for April reports for periods of instabilities that impacted measurements and had a visible impact on sites A/Rs.
    • WLCG Experiments Test Framework (ETF), which replaces the existing SAM/Nagios infrastructure will enter production on 25th of April. ETF has been re-designed and will be deployed in a special configuration and therefore shouldn’t suffer from the same performance issues as those seen with the current SAM/Nagios infrastructure. A summary of the changes impacting sites was presented at the WLCG operations coordination (https://indico.cern.ch/event/466818/). The service change will be transparent in the SAM3 dashboards and starting from 25th of April the source of the measurements in the existing profiles will be switched from the current production to the ETF. More information on ETF can be found in the documentation at http://etf.cern.ch/docs
  • CERN storage services:
    • FTS service upgrade to v 3.4.3, tomorrow 19/04 from 9:30 to 10:30 CET. The upgrade is transparent to the clients.
  • CERN databases:
  • GGUS: (MariaD) Guenter wrote that the instabilities reported by many people in this meeting, ALICE, LHCb, the SCOD and some sites, may be due to 2 obsolete .r0 files found on the web servers and removed. After restarting the web servers the problem seems to be fixed. GGUS:120891 updated.
  • MW Officer:
  • Security:

AOB:

Topic attachments
I Attachment History Action Size Date Who Comment
Unknown file formatpptx GGUS-for-MB-Apr-16.pptx r1 manage 2849.2 K 2016-04-18 - 16:40 MariaDimou GGUS Slide for the 19/4 MB
Edit | Attach | Watch | Print version | History: r23 < r22 < r21 < r20 < r19 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r23 - 2016-04-19 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback